A parent calls a clinic about a child weighing 9.6kg who has a temperature of 102.2F after receiving immunizations. The clinic protocol is to administer acetaminophen 15mg/kg every 4 to 6 hours as needed for fever. The parent has acetaminophen 160mg per 5 ml. How many ml should the nurse instruct the parent to administer for each dose? Amswer in mL.
Answer #1

ANSWER.

* Calculation of amount(ml) of acetaminophen administer for each dose.

           Formula = Desired dose / Dose in hand * Quantity.

      Here

           Ordered dose as per protocol = 15 mg/kg.

            Weight of the child = 9.6 kg.

       Desired dose(D) for a child with 9.6 kg weight = 15 * 9.6

                                                                                 = 144 mg.

                                                    Dose in hand (H) = 160 mg.

                                                    Quantity(Q) = 5 ml.

       As per above formula

              Amount(ml) of drug required per dose = D/H * Q

                                                                            = 144/160 * 5

                                                                            = 4.5 ml.